# The adelic closure of triangle groups

### v1

@@ Abstract.
Motivated by questions arising from billiard trajectories in the regular $n$-gon, McMullen [McM24] defined a pair of functions $\kappa$ and $\delta$ on the cusps $c$ of the corresponding triangle group $\Delta_{n}$ inside $\mathrm{SL}_{2}({\mathcal{O}})$, where ${\mathcal{O}}=\mathbf{Z}[\zeta_{n}+\zeta^{-1}_{n}]$. McMullen asks for which $n$ these functions are congruence, that is, when they only depend on the image of the cusp $c\in\mathbf{P}^{1}(\mathcal{O})$ in $\mathbf{P}^{1}(\mathcal{O}/N)$ for some integer $N$. In this note, we answer McMullen’s questions. We obtain our results by computing the exact closure of $\Delta_{n}\subset\mathrm{SL}_{2}({\mathcal{O}})$ inside $\mathrm{SL}_{2}(\widehat{{\mathcal{O}}})$, where $\widehat{{\mathcal{O}}}$ is the profinite completion of ${\mathcal{O}}$.
The author was supported in part by NSF Grant DMS-2001097.
@@ 1. Introduction
The element $R$ has order $n$, and the quotient $\mathbf{H}/\Delta$ is a hyperbolic orbifold of genus zero with signature $(n/2,\infty,\infty)$ when $n$ is even and $(2,n,\infty)$ when $n$ is odd. If $K=\mathbf{Q}(\zeta+\zeta^{-1})$ and $\mathcal{O}=\mathcal{O}_{K}=\mathbf{Z}[\zeta+\zeta^{-1}]$, then $\Delta\subset\Gamma:=\mathrm{SL}_{2}(\mathcal{O})$. Let $\widehat{\mathcal{O}}$ denote the closure of $\mathcal{O}$ inside the finite adeles $\mathbf{A}^{\kern-0.70004pt{f}}_{K}$. There is a natural map
$\Delta\rightarrow\mathrm{SL}_{2}(\mathcal{O})\hookrightarrow\mathrm{SL}_{2}(\widehat{\mathcal{O}}).$
Let $\widehat{\Delta}$ denote the closure of the image of $\Delta$. (The adelic group $\widehat{\Gamma}:=\mathrm{SL}_{2}(\widehat{\mathcal{O}})$ has a natural topology as a profinite group; we can also view it as the congruence closure of $\mathrm{SL}_{2}(\mathcal{O})$, that is, the profinite group given by the projective limit of the groups $\mathrm{SL}_{2}(\mathcal{O}/d)$ over all integers $d$.) As noted in [McM24, Prop 1.15], the group $\widehat{\Delta}$ has finite index in $\widehat{\Gamma}$. The goal of this note is to describe $\widehat{\Delta}$ as precisely as possible. One application of these results is to answer a number of questions first raised in [McM24]. (The paper [McM24] is itself an offshoot of a number of other papers — including [McM23] and [McM22] — whose goal is to understand billiards in regular polygons.) One of the intentions of [McM24] was to distill a number of purely group theoretical and number theoretical questions in a transparent way, and [McM24] makes no direct reference to billiards. In this paper, we similarly shall not mention these connections, but rather refer the reader (as a starting point) to [McM24, McM23, McM22]. Instead, the questions raised in [McM24] are phrased directly in terms of the arithmetic of the cusps of $\Delta$, and we now recall the basic story of these cusps. The group $\Delta\subset\mathrm{SL}_{2}(\mathbf{R})$ acts on the usual upper half plane $\mathbf{H}$ with finite covolume. The quotient curve (orbifold) $\mathbf{H}/\Delta$ has either one or two cusps ($\infty$ and 00) depending on the parity of $n$; in the usual way one can think about these cusps as certain orbits of $\mathbf{P}^{1}(K)\subset\mathbf{P}^{1}(\mathbf{R})$ under the action of $\Delta$. Very explicitly, if
then $a/c\in\mathbf{P}^{1}(K)$ lies in the orbit of $\infty$ and $b/d$ lies in the orbit of 00. If $[K:\mathbf{Q}]\leq 2$, then every point of $\mathbf{P}^{1}(K)$ is a cusp. If $[K:\mathbf{Q}]>2$, this is no longer true, and there is no known explicit characterization of the cusps. Even when $[K:\mathbf{Q}]=2$ (which occurs for $n=5$, $8$, and $12$), there are still many mysteries regarding the cusps. In these cases, given $x\in\mathbf{P}^{1}(K)$, then there exists a $\gamma\in\Delta$ as in (1) such that either $a/c$ or $b/d$ is equal to $x$. (For more general $n$ this still holds at least if one assumes that $x$ is a cusp of $\Delta$.) But the column $[a,c]$ of $\gamma$ turns out to be unique up to sign, whereas (assuming $[K:\mathbf{Q}]>1$ so the unit group of $K$ is infinite) there are many possible pairs $(a,c)$ for which $a/c=x$. It is precisely the map which sends a cusp $x\in\mathbf{P}^{1}(K)$ to a representative $[a,c]$ (up to sign) with $a/c=x$ which is poorly understood. In [McM24], certain simple functions $\delta$ and $\kappa$ on the set of cusps are considered. For example, when $n$ is even, the map $\kappa$ is the function on the set of cusps $x\in\mathbf{P}^{1}(K)$ which is 00 when $x$ is in the orbit of the cusp $\infty$ and $1$ if $x$ is in the orbit of 00. The definition of $\delta$ is slightly more complicated and is recalled in section 6. McMullen raises the problem as to what extent these maps are congruence, that is, whether they only depend on the image of $x\in\mathbf{P}^{1}(\mathcal{O})$ in $\mathbf{P}^{1}(\mathcal{O}/N)$ for some integer $N$. It is shown in [McM24] that, when the prime factorization of $n$ has certain explicit forms, these maps are congruence. However, it is also shown [McM24, Thm 1.13] that $\delta$ is not congruence when $n=12$. In this paper, we extend this result to show that $\delta$ and $\kappa$ are never congruence invariants in all the remaining cases not addressed in [McM24], generalizing the result (and method) for $\delta$ addressed in [McM24] when $n=12$. In particular, we have (see Theorems 6.1 and 6.3):
Theorem 1.1.
The invariant $\delta$ congruence if and only if either $4\nmid n$ or $n=2^{k}$. The invariant $\kappa$ is congruence unless $n=2m$ where $m$ is odd and has at least two prime factors.
As suggested by the arguments of [McM24], one can establish this result as a consequence of understanding the adelic closure $\widehat{\Delta}$ of $\Delta$ in $\widehat{\Gamma}=\mathrm{SL}_{2}(\widehat{\mathcal{O}})$. We turn to this now, breaking up the result into two parts depending on whether one is considering the prime $p>2$ or $p=2$.
@@
This follows from Lemmas 4.24.6, and 4.8. We now give a precise description of the $2$-adic closure $\widehat{\Delta}_{2}$. In the introduction, we only give the statement in the case when $n$ is not a power of $2$, and leave the statement of the remaining case when $n=2^{k}$ to the main body of the text (see Theorem 5.2). For any (commutative) ring $A$, any ideal $I$, and any group $G\subset\mathrm{SL}_{2}(A)$, let $G(J)$ denote the kernel of the homomorphism $G\rightarrow\mathrm{SL}_{2}(A/J)$. Furthermore, we let $M^{0}(A)$ denote the additive group of $2\times 2$ matrices over $A$ with trace zero. If $A$ is a ring with $2=0$, then $M^{0}(A)$ has a subgroup isomorphic to $A$ consisting of diagonal matrices. The following is proved in Section 3 (note that Theorem 1.3 (3) is [McM24, Thm 1.1]).
@@
Combining these two theorems (together with Theorem 5.2 which explains what happens when $n=2^{k}$), we obtain the following formula for the full index of $\widehat{\Delta}$ in $\widehat{\Gamma}$, namely:
@@ 2. Some preliminaries
In this section, we record some preliminary facts and lemmas which will be used in the sequel. In addition to the field $K=\mathbf{Q}(\zeta+\zeta^{-1})$, we shall also consider $L=\mathbf{Q}(\zeta)$. If we write $n=2^{k}m$ with $m$ odd, then $\mathrm{Gal}(L/\mathbf{Q})\simeq(\mathbf{Z}/n\mathbf{Z})^{\times}=(\mathbf{Z}/m\mathbf{Z})^{\times}\times(\mathbf{Z}/2^{k}\mathbf{Z})^{\times}$, whereas $\mathrm{Gal}(K/\mathbf{Q})$ is the quotient of this group by the image of complex conjugation, given explicitly by $c=(-1,-1)$. If $m>1$, it follows from this description and the natural identification of the inertia group at $2$ of $\mathrm{Gal}(L/\mathbf{Q})$ with $(\mathbf{Z}/2^{k}\mathbf{Z})^{\times}$ that $L/K$ is unramified at all primes above $2$. The ramification index $e$ of (both) fields at the prime $2$ is $e=\varphi(2^{k})$ which equals $2^{k-1}$ if $k\geq 1$ and $1$ otherwise, unless $n=2^{k}$ in which case $e=2^{k-2}$ for $K$. Although $\Delta$ is naturally a subgroup of $\mathrm{SL}_{2}(\mathcal{O}_{K})$, it will sometimes be convenient to work in the larger group $\mathrm{SL}_{2}(\mathcal{O}_{L})$ where certain elements are naturally diagonalizable, making a number of explicit computations easier.
@@ 2.1. The adjoint representation
Let $J$ be an ideal of $\mathcal{O}_{K}$. The ideal $J$ is principal in the ring $\mathcal{O}/J^{2}$ and any choice of generator gives an isomorphism $J/J^{2}\simeq\mathcal{O}/J$. If $\Gamma=\mathrm{SL}_{2}(\mathcal{O}_{K})$, then there is a corresponding identification
@@
A natural approach to understanding the closure of a subgroup $G$ of $\mathrm{SL}_{2}(\mathcal{O}_{\mathfrak{p}})$ is to consider the filtration of $G$ by $G(\mathfrak{p}^{r})$ and understand the graded pieces $G(\mathfrak{p}^{r})/G(\mathfrak{p}^{r+1})$ as modules for the image $G/G(\mathfrak{p})$ of $G$ in $\mathrm{SL}_{2}(\mathcal{O}/\mathfrak{p})$.
@@
2.3. The adjoint representation at $2$
It will be of particular importance when studying the $2$-adic image of $\Delta$ to understand the basic group theory of these modules, which we discuss now. If $n=2^{k}m$ with $m>1$ odd, it is shown in [McM24, Theorem 1.1] that the image of $\Delta$ in $\mathrm{SL}_{2}(\mathcal{O}_{K}/2)$ is isomorphic to $D_{n/2}$ if $k>0$ and $D_{n}$ otherwise. If $\mathfrak{p}$ is a prime of residue characteristic two, then the image of $\Delta$ in $\mathrm{SL}_{2}(\mathcal{O}_{K}/\mathfrak{p})$ is isomorphic to $D_{m}$. (The importance of the image of $\Delta$ in $\mathrm{SL}_{2}(\mathcal{O}_{L}/\mathfrak{p})$ is why we reserve $m$ for the largest odd divisor of $n$, in contrast to [McM24] where $n=m$ or $n=2m$ depending on whether $n$ is odd or not; hopefully this conflict does not cause any confusion.)
@@ 2.3. The adjoint representation at $2$
$T=\left(\begin{matrix}1&2+\zeta+\zeta^{-1}\\ 0&1\end{matrix}\right),\quad U=-\left(\begin{matrix}1&0\\ -1&1\end{matrix}\right),\quad R=U.T.$
The matrix $R$ has order $n$ and eigenvalues $\zeta$ and $\zeta^{-1}$. Moreover, we can explicitly diagonalize $R$ by the matrix of eigenvectors:
@@
Note that this matrix lies in $\mathrm{GL}_{2}(\mathcal{O}_{\mathfrak{p}})$ for any prime $\mathfrak{p}|2$ only if $m>1$ (so that $1+\zeta$ is a $2$-adic unit). In the new basis, we have
@@
The basis (4) will occasionally be more convenient for some computations. Note that it is only defined over the larger field $L=\mathbf{Q}(\zeta)$ rather than $K=\mathbf{Q}(\zeta+\zeta^{-1})$. We only use this basis when $m>1$. Let us now assume that $m>1$. (The case $n=2^{k}$ is considered in Section 5.) Let $D_{m}$ denote the dihedral group of order $2m$. Let $\mathbf{F}$ be the finite field of characteristic $2$ given by adjoining $\zeta+\zeta^{-1}$ to $\mathbf{F}_{2}$ for a primitive $m$th root of unity $\zeta$. (Note that if $\zeta$ is a primitive $2^{k}m$th root of unity, then the image of $\zeta$ in any field of characteristic $2$ is a primitive $m$th root of unity.) The group $D_{m}$ admits a $2$-dimensional representation over $\mathbf{F}$ given by the image of $\Delta$ in $\mathrm{SL}_{2}(\mathcal{O}/\mathfrak{p})$ for a prime $\mathfrak{p}|2$. The action of conjugation makes $M^{0}(\mathbf{F})$ a representation for $D_{m}$. There is a splitting
@@ 2.5. Roots of unity
where $i$ and $j$ are odd contains an element of $\mathcal{O}_{L}/2$ with valuation $12r+$ 1 at $\mathfrak{p}$ and valuation greater than $12r+$ 1 at all other primes above $2$.
@@
Since $\xi-1$ has valuation $1$, the second claim follows from the first (even taking $j=1$). If $\alpha\in\mathcal{O}_{L}$ is expressible as a sum of an even number of roots of unity to odd exponent, then so is
@@
and the valuation of $\alpha(\xi^{2}-1)=\alpha(\xi-1)^{2}\bmod 2$ at any $\mathfrak{p}$ is exactly $2$ plus the valuation of $\alpha$. Hence, by induction, it suffices to consider the case when $r=0$. Certainly there exists an $\eta\in\mathcal{O}$ which is divisible by all primes above $2$ with the exception of a chosen prime $\mathfrak{p}$. Multiplying by the unit $(\zeta-1)$ (which is a unit since $m>1$), we can assume that $\eta$ is expressible as a sum of an even number of roots of unity. Replacing any even odd exponent $\zeta$ by $\zeta^{i+m}$ keeps $\eta$ unchanged modulo $(1-\xi)\mathcal{O}_{L}$, which preserves the property of being co-prime to $\mathfrak{p}$ but not any other prime above $2$, and hence we have constructed the desired element. ∎

### v2

@@ Abstract.
[…]
The author was supported in part by NSF Grants DMS-2001097 and DMS-2450123.
@@ 1. Introduction
The element $R$ has order $n$, and the quotient $\mathbf{H}/\Delta$ is a hyperbolic orbifold of genus zero with signature $(n/2,\infty,\infty)$ when $n$ is even and $(2,n,\infty)$ when $n$ is odd. If $K=\mathbf{Q}(\zeta+\zeta^{-1})$ and $\mathcal{O}=\mathcal{O}_{K}=\mathbf{Z}[\zeta+\zeta^{-1}]$, then $\Delta\subset\Gamma:=\mathrm{SL}_{2}(\mathcal{O})$. Let $\widehat{\mathcal{O}}$ denote the closure of $\mathcal{O}$ inside the finite adeles $\mathbf{A}^{\kern-0.81949pt{f}}_{K}$. There is a natural map
[…]
@@
This follows from Lemmas 4.24.6, and 4.8. Theorem 1.1 determines $\widehat{\Delta}$ up to a precise description of the $2$-adic closure $\widehat{\Delta}_{2}$, which we turn to now. In this introduction, we only give the statement in the case when $n$ is not a power of $2$, and leave the statement of the remaining case when $n=2^{k}$ to the main body of the text (see Theorem 5.2). We first recall some standard definitions:
Definition 1.2.
For any (commutative) ring $A$, any ideal $I$, and any group $G\subset\mathrm{SL}_{2}(A)$, let $G(J)$ denote the kernel of the homomorphism $G\rightarrow\mathrm{SL}_{2}(A/J)$. Furthermore, we let $M^{0}(A)$ denote the additive group of $2\times 2$ matrices over $A$ with trace zero.
If $A$ is a ring with $2=0$, then $M^{0}(A)$ has a subgroup isomorphic to $A$ consisting of diagonal matrices. The following is proved in Section 3 (note that Theorem 1.3 (3) is [McM24, Thm 1.1]).
@@
One application of these results is to answer a number of questions first raised in [McM24]. (The paper [McM24] is itself an offshoot of a number of other papers — including [McM23] and [McM22] — whose goal is to understand billiards in regular polygons.) One of the intentions of [McM24] was to distill a number of purely group theoretical and number theoretical questions in a transparent way, and [McM24] makes no direct reference to billiards. In this paper, we similarly shall not mention these connections, but rather refer the reader (as a starting point) to [McM24, McM23, McM22]. Instead, the questions raised in [McM24] are phrased directly in terms of the arithmetic of the cusps of $\Delta$, and we now recall the basic story of these cusps. The group $\Delta\subset\mathrm{SL}_{2}(\mathbf{R})$ acts on the usual upper half plane $\mathbf{H}$ with finite covolume. The quotient curve (orbifold) $\mathbf{H}/\Delta$ has either one or two cusps ($\infty$ and 0) depending on the parity of $n$; in the usual way one can think about these cusps as certain orbits of $\mathbf{P}^{1}(K)\subset\mathbf{P}^{1}(\mathbf{R})$ under the action of $\Delta$. Very explicitly, if
then $a/c\in\mathbf{P}^{1}(K)$ lies in the orbit of $\infty$ and $b/d$ lies in the orbit of 0. If $K:\mathbf{Q}]\leq 2$, then every point of $\mathbf{P}^{1}(K)$ is a cusp. If $K:\mathbf{Q}]>2$, this is no longer true, and there is no known explicit characterization of the cusps. Even when $K:\mathbf{Q}]=2$ (which occurs for $n=5$, 88, and $12$), there are still many mysteries regarding the cusps. In these cases, given $x\in\mathbf{P}^{1}(K)$, then there exists a $\gamma\in\Delta$ as in (1) such that either $a/c$ or $b/d$ is equal to $x$. (For more general $n$ this still holds at least if one assumes that $x$ is a cusp of $\Delta$.) But the column $[a,c]$ of $\gamma$ turns out to be unique up to sign, whereas (assuming $K:\mathbf{Q}]>1$ so the unit group of $K$ is infinite) there are many possible pairs $(a,c)$ for which $a/c=x$. It is precisely the map which sends a cusp $x\in\mathbf{P}^{1}(K)$ to a representative $[a,c]$ (up to sign) with $a/c=x$ which is poorly understood. In [McM24], certain simple functions $\delta$ and $\kappa$ on the set of cusps are considered. For example, when $n$ is even, the map $\kappa$ is the function on the set of cusps $x\in\mathbf{P}^{1}(K)$ which is 0 when $x$ is in the orbit of the cusp $\infty$ and $1$ if $x$ is in the orbit of 0. The definition of $\delta$ is slightly more complicated and is recalled in section 6. McMullen raises the problem as to what extent these maps are congruence, that is, whether they only depend on the image of $x\in\mathbf{P}^{1}(\mathcal{O})$ in $\mathbf{P}^{1}(\mathcal{O}/N)$ for some integer $N$. It is shown in [McM24] that, when the prime factorization of $n$ has certain explicit forms, these maps are congruence. However, it is also shown [McM24, Thm 1.13] that $\delta$ is not congruence when $n=12$. In this paper, we extend this result to show that $\delta$ and $\kappa$ are never congruence invariants in all the remaining cases not addressed in [McM24], generalizing the result (and method) for $\delta$ addressed in [McM24] when $n=12$. In particular, we have (see Theorems 6.3 and 6.5 and § 6.1):
Theorem 1.4.
The invariant $\kappa$ is congruence unless $n=2m$ where $m$ is odd and has at least two prime factors. The invariant $\delta$ as a function on the cusps equivalent to $\infty$ is congruence if and only if either $4\nmid n$ or $n=2^{k}$. The invariant $\delta$ as a function on all the cusps is congruence if and only if either $n=2^{k}$, $n=2p^{k}$ for an odd prime $p$, or $n$ is odd.
[…]
Remark 1.5.
If $E$ is an elliptic curve over a number field $F$, then, by considering the collection of $p$-adic Galois representations associated to the Tate module $T_{p}(E)$ for each prime $p$, one obtains a representation:
[…]
Combining Theorems 1.1 and 1.3 (together with Theorem 5.2 which explains what happens when $n=2^{k}$), we obtain the following formula for the full index of $\widehat{\Delta}$ in $\widehat{\Gamma}$, namely:
@@ 2. Some preliminaries
In this section, we record some preliminary facts and lemmas which will be used in the sequel. In addition to the field $K=\mathbf{Q}(\zeta+\zeta^{-1})$, we shall also consider $L=\mathbf{Q}(\zeta)$. If we write $n=2^{k}m$ with $m$ odd, then $\mathrm{Gal}(L/\mathbf{Q})\simeq(\mathbf{Z}/n\mathbf{Z})^{\times}=(\mathbf{Z}/m\mathbf{Z})^{\times}\times(\mathbf{Z}/2^{k}\mathbf{Z})^{\times}$, whereas $\mathrm{Gal}(K/\mathbf{Q})$ is the quotient of this group by the image of complex conjugation, given explicitly by c=(1,1)c=(-1,-1). If $m>1$, it follows from this description and the natural identification of the inertia group at $2$ of $\mathrm{Gal}(L/\mathbf{Q})$ with $(\mathbf{Z}/2^{k}\mathbf{Z})^{\times}$ that $L/K$ is unramified at all primes above $2$. The ramification index $e$ of (both) fields at the prime $2$ is $e=\varphi(2^{k})$ which equals $2^{k-1}$ if $k\geq 1$ and $1$ otherwise, unless $n=2^{k}$ in which case $e=2^{k-2}$ for $K$. Although $\Delta$ is naturally a subgroup of $\mathrm{SL}_{2}(\mathcal{O}_{K})$, it will sometimes be convenient to work in the larger group $\mathrm{SL}_{2}(\mathcal{O}_{L})$ where certain elements are naturally diagonalizable, making a number of explicit computations easier.
@@ 2.1. The adjoint representation
Let $J$ be an ideal of $\mathcal{O}_{K}$. The ideal $J$ is principal in the ring $\mathcal{O}/J^{2}$ and any choice of generator gives an isomorphism $J/J^{2}\simeq\mathcal{O}/J$. If $\Gamma=\mathrm{SL}_{2}(\mathcal{O}_{K})$, then (with $\Gamma(J)$ defined as in Definition 1.2) there is a corresponding identification
@@
A natural approach to understanding the closure of a subgroup $G$ of $\mathrm{SL}_{2}(\mathcal{O}_{\mathfrak{p}})$ is to consider the images of $G$ in $\mathrm{SL}_{2}(\mathcal{O}/\mathfrak{p}^{r})$ for each $r$, or equivalently the filtration of $G$ by $G(\mathfrak{p}^{r})$. One then wants to understand the graded pieces $G(\mathfrak{p}^{r})/G(\mathfrak{p}^{r+1})$ as modules for the image $G/G(\mathfrak{p})$ of $G$ in $\mathrm{SL}_{2}(\mathcal{O}/\mathfrak{p})$.
@@
2.3. The adjoint representation at 22
It will be of particular importance when studying the $2$-adic image of $\Delta$ to understand the basic group theory of these modules, which we discuss now. If $n=2^{k}m$ with $m>1$ odd, it is shown in [McM24, Theorem 1.1] that the image of $\Delta$ in $\mathrm{SL}_{2}(\mathcal{O}_{K}/2)$ is isomorphic to $D_{n/2}$ if $k>0$ and $D_{n}$ otherwise. If $\mathfrak{p}$ is a prime of residue characteristic two, then the image of $\Delta$ in $\mathrm{SL}_{2}(\mathcal{O}_{K}/\mathfrak{p})$ is isomorphic to the (typically smaller) group $D_{m}$. (The importance of the image of $\Delta$ in $\mathrm{SL}_{2}(\mathcal{O}_{K}/\mathfrak{p})$ is why we reserve $m$ for the largest odd divisor of $n$, in contrast to [McM24] where $n=m$ or $n=2m$ depending on whether $n$ is odd or not; hopefully this conflict does not cause any confusion.)
@@ 2.3. The adjoint representation at 22
The matrix $R$ has order $n$ and eigenvalues $\zeta$ and $\zeta^{-1}$.
For some purposes, it will be useful to note that we can explicitly diagonalize $R$ by the matrix of eigenvectors:
@@
Note that this matrix lies in $\mathrm{GL}_{2}(\mathcal{O}_{\mathfrak{p}}(\zeta))$ for any prime $\mathfrak{p}|2$ only if $m>1$ (so that $1+\zeta$ is a $2$-adic unit). In the new basis, we have
@@
The basis (4) will occasionally be more convenient for some computations. Note that it is only defined over the larger field $L=\mathbf{Q}(\zeta)$ rather than $K=\mathbf{Q}(\zeta+\zeta^{-1})$. We only use this basis when $m>1$.
Let us now assume that $m>1$. (The case $n=2^{k}$ is considered in Section 5.) Let $D_{m}$ denote the dihedral group of order $2m$. Let $\mathbf{F}$ be the finite field of characteristic $2$ given by adjoining $\zeta+\zeta^{-1}$ to $\mathbf{F}_{2}$ for a primitive $m$th root of unity $\zeta$. (Note that if $\zeta$ is a primitive $2^{k}m$th root of unity, then the image of $\zeta$ in any field of characteristic $2$ is a primitive $m$th root of unity.) There is an isomorphism $\mathcal{O}/\mathfrak{p}=\mathbf{F}$ for any prime $\mathfrak{p}|2$. The group $D_{m}$ admits a $2$-dimensional representation over $\mathbf{F}$ given by the image of $\Delta$ in $\mathrm{SL}_{2}(\mathcal{O}/\mathfrak{p})$ for a prime $\mathfrak{p}|2$. The action of conjugation makes $M^{0}(\mathbf{F})$ a representation for $D_{m}$. There is a splitting
@@ 2.5. Roots of unity
where $i$ and $j$ are odd contains an element of $\mathcal{O}_{L}/2$ with valuation $2r+1$ at $\mathfrak{p}$ and valuation greater than $2r+1$ at all other primes above $2$.
@@
Since $\xi-1$ has valuation $1$ at all primes above $2$, the second claim follows from the first (even taking $j=1$). If $\alpha\in\mathcal{O}_{L}$ is expressible as a sum of an even number of roots of unity to odd exponent, then so is
@@
and the valuation of $\alpha(\xi^{2}-1)=\alpha(\xi-1)^{2}\bmod 2$ at any $\mathfrak{p}$ is exactly $2$ plus the valuation of $\alpha$. Hence, by induction, it suffices to consider the case when $r=0$. Certainly there exists an $\eta\in\mathcal{O}$ which is divisible by all primes above $2$ with the exception of a chosen prime $\mathfrak{p}$. Multiplying by the unit $(\zeta-1)$ (which is a unit since $m>1$), we can assume that $\eta$ is expressible as a sum of an even number of roots of unity. Replacing any even odd exponent $\zeta$ by $\zeta^{i+m}$ keeps $\eta$ unchanged modulo $(1-\xi)\mathcal{O}_{L}=\prod\mathfrak{p}_{i}$, which preserves the property of being co-prime to $\mathfrak{p}$ but not having positive valuation at every other prime above $2$, and hence we have constructed the desired element. ∎